OpenStudy (anonymous):

If you invest $10,000 at 5% interest compounded continuously, what is the average amount in your account over one year?

OpenStudy (anonymous):

5% interest compounded continuously. What is meant by continuously? Is that what it says in your question?

OpenStudy (anonymous):

yes that is what my question says

OpenStudy (anonymous):

Ah I see.

OpenStudy (anonymous):

You take the limit of n approaching to infinity.

OpenStudy (kropot72):

\[A=Pe ^{rt}\] \[\int\limits_{}^{}A\ dt=\frac{Pe ^{rt}}{r}+C\] \[\int\limits_{t=0}^{t=1}A\ dt=\frac{1}{0.05}(10000e ^{0.05}-10000)\] Can you calculate the average value now?

OpenStudy (anonymous):

yes thank you

OpenStudy (kropot72):

Cool!